Subject: 7/16 Users Committee agenda, telescope report

From: Bruce Gillespie

Submitted: Thu, 12 Jul 2007 14:28:34 -0600

Message number: 1041 (previous: 1040, next: 1042 up: Index)


Suzanne asked me to send you the agenda, the call-in information, and  
various informational reports for the Monday 7/16 3.5-m Users  
Committee telecon, which are given below.  Committee members, please  
be prepared to address the various agenda items from your users' and  
institution's perspectives.

The call-in details are:


number: 866-719-0837
passcode: 1187123

time & date: 8:30am PDT.  Monday, 16 July 2007


   o User feedback and comments - Committee members

   o Discussion of telescope/instruments report - Klaene

   o New SPIcam/TUI interface - Klaene

   o Other topics - all

   o Actions from earlier meetings - group

Actions from earlier meetings - group

[open from previous months]:


[new actions from last month]:

ACTION:  Jon Holtzman will investigate and report on the tip-tilt  
system at the TNG telescope on La Palma.


  3.5-m Telescope, Instruments, and CIF Projects Highlights, 6/10/07  
through 7/5/07

                                  Mark Klaene

0) Overview

Monsoonal weather started a little early this year with moderate rain  
and evening clouds limiting observations for much of the past 2  
weeks. The June engineering time was clouded out; however, some in- 
dome testing of the SPIcam TUI interface and further training of Jim  
Davenport made the time useful.  Bill Ketzeback has returned from his  
family medical leave.

1) Telescope

No significant problems this period.  Tracking tuning was not  
accomplished, but tracking has generally been acceptable, not  
requiring that we pre-empt observing.

The summer shutdown is still on track for August 20 to September 16.   
The driver is the primary mirror re-aluminization.  We have a limited  
number of days on the sky for recovery; hopefully the weather will  
cooperate and allow us to return to science on schedule.

2) Instruments

Agile is preparing for more commissioning runs at the telescope this  
month.  The fiber interface was working when it left Princeton  
Instruments.  The focal reducer is planned to be completed and ready  
for commissioning during July.

The replacement DIS Red high-resolution grating has been received at  
JHU for testing.  Delivery to the observatory and installation in DIS  
is expected later this month.

The Lakeshore temperature controller failed on both DIS cameras-- 
repair is in work.  This leaves us with no DIS red heater and only  
the ion pump controllers as a way to gauge temp and vacuum.

We replaced the ThAr lamp on the Echelle.

Incorporation of SPIcam into TUI has been released to users.  Some  
scripts are still under development.

NA2 guider camera had an intermittent network connection failure  
develop.  The cause was traced to a broken network cable.  A more  
flexible cable has been installed.

NIC-FPS is scheduled to be taken off line and sent to CU on the 26th  
of July.  Work will include implementing 16-channel readout  
capability and F-P Etalon installation and testing.

3) Engineering and CIF projects


Rotator upgrades: two machining work orders have been let and the PO  
for the motor and accessories has been signed.  Expect delivery late  

Axis controllers: bench setup has been built and is undergoing  
testing of the new axis controllers.


Tertiary mounting:  discussions are ongoing as to the need for tip/ 
tilt  to be incorporated.  Input from the scientific community is  
required to determine if fast guiding at the tertiary is desired.  If  
interested please let Larry Carey know (

Direct drives: research into direct drives for the azimuth and  
altitude drives continues.

4) Miscellaneous

Jim Davenport completed his training as a summer fill-in Observing  
Specialist.  Jim will be soloing during the rest of July and August.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1041 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to